Lithium has two known stable isotopes- lithium-6 and lithium-7. Lithium-6 has an isotopic mass of 6.015 amu and percent abundance of 7.59%; lithium-7 has a mass of 7.016 amu at 92.41 % abundance. Calculate the average atomic mass of lithium.

1
Expert's answer
2022-02-10T12:39:01-0500

For isotope 1 (lithium-6), atomic mass = "6.015\times 7.59/100=0.4565385"

For lithium-7, atomic mass"=7.016\times 92.41/100=6.4834856"

Average atomic mass of lithium"=0.4565385+6.4834856=6.9400241 amu"